I re-submitted this question. I had missing data, sorry!

An A-frame lake house is 21 feet wide.If the roof of the home makes a 45 degree angle with the base of the home,what is the length of the house from ground level to the peak of the roof?
Larson, R., Hostetler, R. P., & Edwards, B. H. (2005). Trigonometry a graphing approach (4th ed.). Boston: Houghton Mifflin.

Answer by checkley77(12569)

(180-45)/2=135/2=67.5 is the angle to the peak.
tan67.5=2.414
2.414(21/2)=25.34 is the height of the A-frame
10.5^2+25.34^2=c^2
110.25+642.12=c^2
752.366=c^2
c=sqrt752.366
c=27.43 ft. is the distance from the ground to the peak.